OpenStudy (anonymous):

Solve each system by graphing y=-1/3x+3 y=-2x-2

OpenStudy (anonymous):

this is simple

OpenStudy (anonymous):

-1/3*-3=1+3=4 so (-3,4) (-3,4) where are graph choices?

OpenStudy (anonymous):

just write it with a regular graph

OpenStudy (anonymous):

ok are those plots on the graph?

OpenStudy (anonymous):

yeah I am sure they are

OpenStudy (anonymous):

ok that's your answer